# Visual Odometry with Transformers

### v1

@@ 4 Experiments
Datasets. Our primary training leverages ARKitScenes (Baruch et al., 2021), a large-scale dataset, providing rich and diverse indoor scenes. We further incorporate ScanNet (Dai et al., 2017) to broaden indoor coverage and KITTI (Geiger et al., 2012) to expose the model to outdoor pose distributions. Out-of-distribution performance is evaluated on TUM_RGBD (Sturm et al., 2012) which is not included in the training set.
@@
ARKitScenes (Baruch et al., 2021) is a large-scale indoor dataset collected with LiDAR-equipped mobile devices, comprising over 5,000 sequences from 1,661 environments and totaling more than 11 million frames, making it one of the most comprehensive resources for indoor 3D scene understanding. To complement it, ScanNet (Dai et al., 2017) provides over 2 million frames across 1,613 scenes. For outdoor settings, KITTI (Geiger et al., 2012) offers RGB images and ground-truth poses collected with GPS and LIDAR. Finally, TUM_RGBD (Sturm et al., 2012) features indoor recordings with motion-capture ground truth. The datasets exhibit variation in camera calibration, poses, image resolution, as well as real-world artifacts such as rolling shutter, motion blur, and lighting variation.
A note on alignment. While most existing methods report evaluation metrics after applying rigid alignment and scale correction to the predicted trajectory, we believe this practice can be misleading for real-world applications, where ground-truth trajectories are unavailable for alignment. Although early approaches relied on this convention due to the inability to robustly estimate the scale of a scene, the assumption is less justified in the context of modern large-scale 3D models. Therefore, all experimental results are based on unaligned predictions.
@@
We evaluate visual odometry using Absolute Translation Error (ATE) and Absolute Rotation Error (ARE) (Teed et al., 2023; Chen et al., 2024). ATE is the RMSE of translation between estimated and ground-truth trajectories, while ARE is the RMSE of angular orientation differences. In addition, we report Relative Translation Error (RTE) and Relative Rotation Error (RRE), which measure the drift in translation and rotation over unit trajectory segments. In the tables, results are color-coded to indicate ranking: best, second-best, and third-best.
@@
Baselines. We compare our method against state-of-the-art visual odometry models (Teed et al., 2023; Chen et al., 2024). For DPVO (Teed et al., 2023), results are averaged over three runs with different random seeds to account for variability. We further benchmark against recent large-scale 3D models that directly predict camera poses (Wang et al., 2025a; Wang & Agapito, 2025). Because these models cannot process long video sequences in a single pass, inputs are split into temporally continuous chunks, with maximum lengths of 30 and 90 frames on our GPUs, respectively; the predicted poses are then concatenated to form the full trajectory. We also compare with a classical (Campos et al., 2021) and a state-of-the-art (Murai et al., 2025) monocular SLAM systems, both using only RGB input. For fairness, loop closure is disabled during evaluation. Finally, we include a recent end-to-end visual odometry model (Françani & Maximo, 2025) in our comparisons.
@@
Implementation details. We use a frozen CroCo (Weinzaepfel et al., 2022) backbone trained within the DUST3R (Wang et al., 2024) framework, consisting of 300 million parameters. We employ 12 alternating time-space attention blocks, totalling 200 million parameters, for the decoder. Our model is trained with the AdamW (Loshchilov & Hutter, 2019) optimizer for 300 epochs. We adopt a cosine learning rate schedule with an initial learning rate of 0.00001 and a warmup phase of 30 epochs. Our model takes 8 input views, sampled at intervals of 3 frames. Input frames are resized to a resolution of $\times224$. Training runs on 12 NVIDIA RTX H100 GPUs for 5 days.
@@ 4.1 Visual Odometry Results
Method w/o Bundle w/o Camera ATE ARE RTE RRE
Adjustment Parameters [m] $\downarrow$ [deg] $\downarrow$ [m] $\downarrow$ [deg] $\downarrow$
ORB-SLAM3 (Campos et al., 2021) 2.58 111.58 0.17 4.01
DPVO (Teed et al., 2023) 5.48 18.68 0.04 0.48
LeapVO (Chen et al., 2024) 28.31 49.25 0.62 1.09
TSFormer (Françani & Maximo, 2025) 285.22 97.53 0.52 0.62
CUT3R (Wang & Agapito, 2025) 2.42 97.70 0.12 4.86
VGGT (Wang et al., 2025a) 2.94 92.37 0.15 4.81
Mast3r-SLAM-VO (Murai et al., 2025) 0.60 14.40 0.36 17.90
VoT (Ours) 0.51 13.40 0.01 0.44
Methods are evaluated using all frames of the test split without alignment to the ground-truth.
Table 1: Pose estimation metrics on ARKitScenes indoor settings. VoT shows superior performance across all metrics without relying on post-optimization and camera parameters.
VoT is competitive without bells and whistles. As shown in Tabs.˜1, 2 and 3, VoT achieves competitive performance on both indoor and outdoor datasets, despite not relying on hand-crafted components. This is particularly notable in real-world settings, where unaligned trajectory metrics are more realistic due to the lack of ground-truth poses.
Large 3D models such as CUT3R (Wang & Agapito, 2025) and VGGT (Wang et al., 2025a), although trained on diverse multi-task 3D datasets - including both ScanNet and ARKitScenes - fail to generalize effectively. CUT3R and VGGT exhibit substantial drift in long sequences, while Mast3r-SLAM-VO (Murai et al., 2025) suffers from scale ambiguity and sparse predictions, resulting in degraded ATE and failure on the KITTI dataset.
Among odometry-specific methods, VoT exhibits strong generalization across datasets. In contrast, LeapVO (Chen et al., 2024), trained solely on synthetic data, struggles on real-world videos. DPVO (Teed et al., 2023) achieves competitive results on ScanNet but deteriorates on ARKitScenes and KITTI, particularly in scale estimation (ATE), limiting its applicability in the wild. ORB-SLAM3 (Campos et al., 2021) fails to reliably estimate scale, leading to poor performance even with loop closure and bundle adjustment enabled. TSFormer (Françani & Maximo, 2025), trained on a small dataset (i.e., KITTI (Geiger et al., 2012)), without strong encoders or robust rotation representations, fails to predict accurate trajectories.
Method w/o Bundle w/o Camera ATE ARE RTE RRE
Adjustment Parameters [m] $\downarrow$ [deg] $\downarrow$ [m] $\downarrow$ [deg] $\downarrow$
ORB-SLAM3 (Campos et al., 2021) 1.91 100.42 0.04 3.75
LeapVO (Chen et al., 2024) 10.84 43.60 0.09 0.43
DPVO (Teed et al., 2023) 1.75 5.91 0.02 0.34
TSFormer (Françani & Maximo, 2025) 285.22 97.53 0.52 0.62
CUT3R (Wang & Agapito, 2025) 4.85 135.52 0.02 0.65
VGGT (Wang et al., 2025a) 1.56 31.73 0.06 1.42
Mast3r-SLAM-VO (Murai et al., 2025) 0.99 8.40 0.58 6.96
@@
Methods are evaluated using all frames of the test split without alignment to the ground-truth.
Table 2: Pose estimation metrics on ScanNet indoor settings. VoT shows competitive performance without relying on bundle adjustment or requiring camera parameters. Moreover, our method accurately predicts absolute poses, as demonstrated by the low ATE.
Method w/o Bundle w/o Camera ATE ARE RTE RRE
Adjustment Parameters [m] $\downarrow$ [deg] $\downarrow$ [m] $\downarrow$ [deg] $\downarrow$
ORB-SLAM3 (Campos et al., 2021) 217.06 51.23 1.05 1.20
DPVO (Teed et al., 2023) 194.55 0.94 0.85 0.06
LeapVO (Chen et al., 2024) 211.80 39.78 0.94 0.96
TSFormer (Françani & Maximo, 2025) 82.05 22.84 0.23 0.24
CUT3R (Wang & Agapito, 2025) 112.36 22.07 0.71 0.58
VGGT (Wang et al., 2025a) 205.67 16.32 2.13 0.25
Mast3r-SLAM-VO (Murai et al., 2025)
@@
Methods are evaluated using all frames of the test split without alignment to the ground-truth.
Table 3: Pose estimation metrics on KITTI outdoor settings. VoT is capable of modeling very different outdoor camera pose distribution, showing competitive performance compared to counterparts. "–" indicates the failure of a method.
Method w/o Bundle w/o Camera ATE ARE RTE RRE
Adjustment Parameters [m] $\downarrow$ [deg] $\downarrow$ [m] $\downarrow$ [deg] $\downarrow$
ORB-SLAM3 (Campos et al., 2021) 1.65 76.14 3.16 0.03
DPVO (Teed et al., 2023) 0.94 3.02 0.07 0.44
LeapVO (Chen et al., 2024) 1.06 7.36 0.03 0.56
TSFormer (Françani & Maximo, 2025) 135.49 84.29 0.60 2.78
CUT3R (Wang & Agapito, 2025) 0.87 29.40 0.14 1.41
VGGT (Wang et al., 2025a) 1.77 31.66 0.08 1.82
Mast3r-SLAM-VO (Murai et al., 2025) 2.31 145.22 0.08 2.02
@@
Methods are evaluated using all frames of the test split without alignment to the ground-truth.
Table 4: Pose estimation metrics on TUM_RGBD out-of-distribution setting. While optimization-based methods shows strong performance in terms of ARE, VoT is capable of generalizing to unseen data and camera parameters and shows superior performance in ATE and RTE.
@@ 4.2 Ablation Studies
We conduct ablation studies to assess key design choices and analyze scaling trends with respect to model capacity and dataset size. For architectural ablations (e.g., attention mechanisms, backbone, and rotation representation), we use two input views with a stride of 10 and train for 150 epochs. For scaling experiments (e.g., varying dataset size), models are trained to convergence for 300 epochs to accurately reflect scalability. All ablations are evaluated on the ScanNet (Dai et al., 2017) dataset.
VoT generalizes to unseen datasets and camera parameters. Tab.˜4 illustrates how VoT performs on the dataset not seen during training. Moreover, the camera parameters on the dataset are drastically different from the ones in the training set.
@@
VoT exhibits robust scaling behavior. Fig.˜3 illustrates how VoT scales with respect to training data size and model capacity. VoT shows consistent performance improvements with both increased training data and a larger number of trainable parameters. Due to computational limits, we cap the number of decoder layers at 12. We analyze the scaling behavior with respect to encoder size and number of input views in the Supplementary Material.
Method ATE[m] $\downarrow$ ARE[deg] $\downarrow$ RTE[m] $\downarrow$ RRE[deg] $\downarrow$
@@
Table 5: Backbone ablation. * denotes the model trained within the Dust3r (Wang et al., 2024) framework with 3D supervision. The results highlight the critical role of training data in determining feature extractor performance, with geometry-focused 3D supervision yielding better results across all metrics.
Pre-training data is the key determinant of backbone performance. We compare features extracted from various pre-trained backbones in Tab.˜5. While DINOv2 (Oquab et al., 2024) and CroCoV2 (Weinzaepfel et al., 2022) yield comparable results, the CroCoV2 encoder trained within the Dust3R (Wang et al., 2024) framework that uses geometry-focused 3D data - achieves substantially better performance (i.e., approximately $2\times$ lower error across all metrics). These results suggest that the quality and relevance of pre-training data, rather than backbone architecture, are critical for learning effective features in visual odometry.
@@
Time-space attention achieves superior performance with lower computational cost. Tab.˜6 compares full self-attention and time-space attention. Time-space attention not only reduces computational cost but also consistently improves accuracy. We attribute this to its matching-like spatial mechanism (see Fig.˜2), where patches at corresponding locations across frames exchange temporal information, enhancing motion-aware feature encoding.
Rotation Rep. ATE[m] $\downarrow$ ARE[deg] $\downarrow$ RTE[m] $\downarrow$ RRE[deg] $\downarrow$
@@
Plucker Rays (Zhang et al., 2024) 0.66 16.11 0.05 1.48
@@
SO(3) projection enhances rotation accuracy. In Tab.˜7, we evaluate different rotation representations. Projecting outputs onto the nearest valid rotation matrix on the SO(3) manifold using the Frobenius norm consistently yields the best results, highlighting its suitability for our formulation.

### v2

@@ 4 Experiments
Datasets. Our primary training leverages ARKitScenes [baruch2022arkit], a large-scale dataset, providing rich and diverse indoor scenes. We further incorporate ScanNet [dai2017scannet] to broaden indoor coverage and KITTI [geiger2012kitti] to expose the model to outdoor pose distributions. Out-of-distribution performance is evaluated on TUM_RGBD [sturm2012tumrgbd] which is not included in the training set. ARKitScenes [baruch2022arkit] is a large-scale indoor dataset collected with LiDAR-equipped mobile devices, comprising over 5,000 sequences from 1,661 environments and totaling more than 11 million frames, making it one of the most comprehensive resources for indoor 3D scene understanding. To complement it, ScanNet [dai2017scannet] provides over 2 million frames across 1,613 scenes. For outdoor settings, KITTI [geiger2012kitti] offers RGB images and ground-truth poses collected with GPS and LIDAR. Finally, TUM_RGBD [sturm2012tumrgbd] features indoor recordings with motion-capture ground truth. The datasets exhibit variation in camera calibration, poses, image resolution, as well as real-world artifacts such as rolling shutter, motion blur, and lighting variation.
@@
Metrics. We evaluate visual odometry using Absolute Translation Error (ATE) and Absolute Rotation Error (ARE) [teed2023dpvo, chen2024leapvo]. ATE is the RMSE of translation between estimated and ground-truth trajectories, while ARE is the RMSE of angular orientation differences. In addition, we report Relative Translation Error (RTE) and Relative Rotation Error (RRE), which measure the drift in translation and rotation over unit trajectory segments. In the tables, results are coded best and $\underline{\text{second best}}$.
A note on alignment. While it is common practice to report evaluation metrics after applying rigid alignment and scale correction to the predicted trajectory, we believe this practice can be impractical for real-world applications, where ground-truth is unavailable. To take this real-world constrain into account we report both aligned and unaligned metrics.
Figure 4: Visualized Predicted Trajectories from Test Splits of ScanNet and ARKit Datasets. Most methods cannot predict a coherent trajectory without aligning to the ground truth. In contrast, VoT predicts an accurate trajectory without alignment.
@@
Baselines. We compare our method against state-of-the-art visual odometry models [teed2023dpvo, chen2024leapvo]. For DPVO [teed2023dpvo], results are averaged over three runs with different random seeds to account for variability. We further benchmark against recent large-scale 3D models that directly predict camera poses [wang2025vggt, wang20243d_reconstruction_with_spatial_memory]. Because these models cannot process long video sequences in a single pass, inputs are split into temporally continuous chunks, with maximum lengths of 30 and 90 frames on our GPUs, respectively; the predicted poses are then concatenated to form the full trajectory. We also compare with a classical [campos2021orbslam3] and a state-of-the-art [murai2024mast3rslam] monocular SLAM systems, both using only RGB input. For fairness, loop closure is disabled during evaluation. Finally, we include a recent end-to-end visual odometry model [francani2025tsformer] in our comparisons.
Method ATE ARE RTE RRE
[m] $\downarrow$ [deg] $\downarrow$ [m] $\downarrow$ [deg] $\downarrow$
Aligned to GT
ORB-SLAM3 [campos2021orbslam3]†∗ 1.08 133.92 0.03 4.01
DPVO [teed2023dpvo]†∗ 0.49 20.01 0.02 0.49
LeapVO [chen2024leapvo]†∗ 0.89 58.42 0.04 1.10
TSFormer [francani2025tsformer] 141.34 56.32 0.37 0.62
CUT3R [wang20243d_reconstruction_with_spatial_memory] 0.67 82.61 0.09 4.86
VGGT [wang2025vggt] 2.26 124.64 0.16 4.81
Mast3r-SLAM-VO [murai2024mast3rslam] 0.28 25.49 0.19 17.89
VoT (Ours) 0.25 10.89 0.01 0.44
Unaligned to GT
ORB-SLAM3 [campos2021orbslam3]†∗ 2.58 111.58 0.17 4.01
DPVO [teed2023dpvo]†∗ 5.48 18.68 0.04 0.48
LeapVO [chen2024leapvo]†∗ 28.31 49.25 0.62 1.09
TSFormer [francani2025tsformer] 285.22 97.53 0.52 0.62
CUT3R [wang20243d_reconstruction_with_spatial_memory] 2.42 97.70 0.12 4.86
VGGT [wang2025vggt] 2.94 92.37 0.15 4.81
Mast3r-SLAM-VO [murai2024mast3rslam] 0.60 14.40 0.36 17.90
VoT (Ours) 0.51 13.40 0.01 0.44
Table 1: Pose estimation metrics on ARKitScenes indoor settings. Methods marked with use bundle adjustment and methods marked with use camera parameters. Provided enough training data, VoT shows superior performance.
@@
Implementation details. We use a frozen CroCo [philippe2022croco] backbone trained within the DUST3R [wang2024dust3r] framework, consisting of 300 million parameters. We employ 12 alternating time-space attention blocks, totalling 200 million parameters, for the decoder. Our model is trained with the AdamW [loshchilov2017adamw] optimizer for 300 epochs. We adopt a cosine learning rate schedule with an initial learning rate of 0.00001 and a warmup phase of 30 epochs. Our model takes 8 input views, sampled at intervals of 3 frames. Input frames are resized to a resolution of $\times224$. Training runs on 12 NVIDIA RTX H100 GPUs for 5 days.
@@ 4.1 Visual Odometry Results
VoT is competitive without bells and whistles. As shown in Tabs. 1, 2 and 3, VoT achieves competitive performance on both indoor and outdoor datasets, despite not relying on hand-crafted components. This is particularly notable in real-world settings, where unaligned trajectory metrics are more realistic due to the lack of ground-truth poses.
Large 3D models such as CUT3R [wang20243d_reconstruction_with_spatial_memory] and VGGT [wang2025vggt], although trained on diverse multi-task 3D datasets - including both ScanNet and ARKitScenes - fail to generalize effectively. CUT3R and VGGT exhibit substantial drift in long sequences, while Mast3r-SLAM-VO [murai2024mast3rslam] suffers from scale ambiguity and sparse predictions, resulting in degraded ATE and failure on the KITTI dataset.
Among odometry-specific methods, VoT exhibits strong generalization across datasets. In contrast, LeapVO [chen2024leapvo], trained solely on synthetic data, struggles on real-world videos. DPVO [teed2023dpvo] achieves competitive results on ScanNet and KITTI but fails to accurately estimate the scale (ATE). ORB-SLAM3 [campos2021orbslam3] fails to reliably estimate scale, leading to poor performance even with loop closure and bundle adjustment enabled. TSFormer [francani2025tsformer], trained on a small dataset, without strong encoders or robust rotation representations, fails to predict accurate trajectories.
Method ATE ARE RTE RRE
[m] $\downarrow$ [deg] $\downarrow$ [m] $\downarrow$ [deg] $\downarrow$
Aligned to GT
ORB-SLAM3 [campos2021orbslam3]†∗ 1.07 141.78 0.04 3.75
DPVO [teed2023dpvo]†∗ 0.16 8.58 0.02 0.34
LeapVO [chen2024leapvo]†∗ 0.74 72.78 0.02 0.43
TSFormer [francani2025tsformer] 105.21 61.42 0.32 0.45
CUT3R [wang20243d_reconstruction_with_spatial_memory] 0.38 24.13 0.02 0.65
VGGT [wang2025vggt] 1.08 80.84 0.07 1.42
Mast3r-SLAM-VO [murai2024mast3rslam] 0.22 19.31 0.22 6.59
VoT (Ours) 0.15 8.50 0.01 0.42
Unaligned to GT
ORB-SLAM3 [campos2021orbslam3]†∗ 1.91 100.42 0.04 3.75
DPVO [teed2023dpvo]†∗ 1.75 5.91 0.02 0.34
LeapVO [chen2024leapvo]†∗ 10.84 43.60 0.09 0.43
TSFormer [francani2025tsformer] 285.22 97.53 0.52 0.62
CUT3R [wang20243d_reconstruction_with_spatial_memory] 4.85 135.52 0.02 0.65
VGGT [wang2025vggt] 1.56 31.73 0.06 1.42
Mast3r-SLAM-VO [murai2024mast3rslam] 0.99 8.40 0.58 6.96
@@
Table 2: Pose estimation metrics on ScanNet indoor settings. Methods marked with use bundle adjustment and methods marked with use camera parameters. Without bells and whistles, VoT shows competitive performance.
Method ATE ARE RTE RRE
[m] $\downarrow$ [deg] $\downarrow$ [m] $\downarrow$ [deg] $\downarrow$
Aligned to GT
ORB-SLAM3 [campos2021orbslam3]†∗ 80.81 157.31 2.82 1.29
DPVO [teed2023dpvo]†∗ 9.74 13.63 0.93 0.06
LeapVO [chen2024leapvo]†∗ 48.86 67.98 0.39 0.96
TSFormer [francani2025tsformer] 15.77 93.57 0.45 0.74
CUT3R [wang20243d_reconstruction_with_spatial_memory] 17.92 73.76 0.50 0.58
VGGT [wang2025vggt] 45.07 98.23 1.78 0.51
Mast3r-SLAM-VO [murai2024mast3rslam]
VoT (Ours) 9.66 31.12 0.32 0.42
Unaligned to GT
ORB-SLAM3 [campos2021orbslam3]†∗ 217.06 51.23 1.05 1.29
DPVO [teed2023dpvo]†∗ 194.55 0.94 0.85 0.06
LeapVO [chen2024leapvo]†∗ 211.80 39.78 0.94 0.96
TSFormer [francani2025tsformer] 140.10 28.45 0.78 0.74
CUT3R [wang20243d_reconstruction_with_spatial_memory] 112.36 22.07 0.71 0.58
VGGT [wang2025vggt] 205.67 16.32 2.13 0.51
Mast3r-SLAM-VO [murai2024mast3rslam]
@@
Table 3: Pose estimation metrics on KITTI outdoor settings. Methods marked with use bundle adjustment and methods marked with use camera parameters. ✗ denotes the methods that failed on some of the sequences. Even though KITTI training data amounts to only several videos, VoT shows competitive performance.
Method ATE ARE RTE RRE
[m] $\downarrow$ [deg] $\downarrow$ [m] $\downarrow$ [deg] $\downarrow$
Aligned to GT
ORB-SLAM3 [campos2021orbslam3]†∗ 0.80 147.39 0.05 3.16
DPVO [teed2023dpvo]†∗ 0.10 3.08 0.04 0.44
LeapVO [chen2024leapvo]†∗
TSFormer [francani2025tsformer] 0.82 153.35 0.05 0.92
CUT3R [wang20243d_reconstruction_with_spatial_memory] 0.87 29.42 0.02 1.41
VGGT [wang2025vggt] 1.05 70.34 0.03 0.77
Mast3r-SLAM-VO [murai2024mast3rslam] 0.83 150.51 1.36 60.03
VoT (Ours) 0.21 7.91 0.01 0.65
Unaligned to GT
ORB-SLAM3 [campos2021orbslam3]†∗ 1.65 76.14 0.03 3.16
DPVO [teed2023dpvo]†∗ 0.94 3.01 0.07 0.44
LeapVO [chen2024leapvo]†∗
TSFormer [francani2025tsformer] 200.87 105.86 0.58 0.92
CUT3R [wang20243d_reconstruction_with_spatial_memory] 0.51 60.86 0.02 1.41
VGGT [wang2025vggt] 1.61 32.66 0.03 0.77
Mast3r-SLAM-VO [murai2024mast3rslam] 1.21 47.81 1.32 60.03
@@
Table 4: Pose estimation metrics on TUM_RGBD out-of-distribution setting. Methods marked with use bundle adjustment and methods marked with use camera parameters. ✗denotes the methods that failed on some of the sequences. VoT performs well on unaligned metrics but lags behind in aligned settings.
@@ 4.2 Ablation Studies
We conduct ablation studies to assess key design choices and analyze scaling trends with respect to model capacity and dataset size. For architectural ablations (e.g., attention mechanisms, backbone, and rotation representation), we use two input views with a stride of 10 and train for 150 epochs. For scaling experiments (e.g., varying dataset size), models are trained to convergence for 300 epochs to reflect scalability accurately. All ablations are evaluated on the ScanNet [dai2017scannet] dataset.
VoT delivers real-time prediction. To compare the speed between approaches, we measure their runtime on ScanNet [dai2017scannet] scene0000_00 using the same environment. As shown in Fig. 1, VoT achieves a considerably faster runtime, more than $3\times$ speedup, compared to existing methods. This efficiency stems from its compact, end-to-end architecture, in contrast to larger 3D models. Moreover, unlike LeapVO, DPVO, and Mast3R-SLAM-VO, VoT does not rely on additional processing steps or bundle adjustment, contributing to both its speed and strong performance.
VoT is competitive on unseen datasets and camera parameters. Tab. 4 illustrates how VoT performs on a dataset not seen during training. Moreover, the camera parameters of the dataset are drastically different from those in the training set. While DPVO leverages effective bundle-adjustment for superior performance, VoT demonstrates competitive results and robust generalization compared to other methods.
@@
VoT exhibits robust scaling behavior. Fig. 5 illustrates how VoT scales with respect to training data size and model capacity. VoT shows consistent performance improvements with both increased training data and a larger number of trainable parameters. Due to computational limits, we cap the number of decoder layers at 12. We analyze the scaling behavior with respect to encoder size and number of input views in the Supplementary Material.
Method ATE ARE RTE RRE
[m] $\downarrow$ [deg] $\downarrow$ [m] $\downarrow$ [deg] $\downarrow$
@@
Table 5: Backbone ablation. CroCoV2* denotes the model trained within the Dust3r [wang2024dust3r] framework, and DinoV2* denotes the backbone train in VGGT framework [wang2025vggt]. The results highlight the critical role of the architecture of the backbone, with geometry-focused supervision yielding better results across all metrics.
@@
Method ATE ARE RTE RRE
[m] $\downarrow$ [deg] $\downarrow$ [m] $\downarrow$ [deg] $\downarrow$
@@
Plucker Rays [zhang2024raydiffusion] 0.66 16.11 0.05 1.48
@@
Architecture is the key determinant of backbone performance. We compare features extracted from various pre-trained backbones in Tab. 5. While DINOv2 [oquab2024dinov2] and CroCoV2 [philippe2022croco] yield comparable results, the CroCoV2 encoder trained within the Dust3R [wang2024dust3r] framework that uses a geometry-focused architectur- achieves substantially better performance (i.e., approximately $2\times$ lower error across all metrics). Moreover, CroCoV2 trained within Dust3R framework shows superior performance to DinoV2 backbone trained on comparable amounts of data [wang2025vggt] and similar tasks. This suggests that the backbone architecture is the key components for a strong performance for our method.
SO(3) projection enhances rotation accuracy. In Tab. 7, we evaluate different rotation representations. Projecting outputs onto the nearest valid rotation matrix on the SO(3) manifold using the Frobenius norm consistently yields the best results, highlighting its suitability for our formulation.
Time-space attention achieves superior performance with lower computational cost. Tab. 6 compares full self-attention and time-space attention. Time-space attention not only reduces computational cost but also consistently improves accuracy. We attribute this to its matching-like spatial mechanism (see Fig. 3), where patches at corresponding locations across frames exchange temporal information, enhancing motion-aware feature encoding.
Limitations and future work. We do not claim VoT can generalize to all the datasets. Since it is trained on mostly static environments, performance may be limited in dynamic settings. Future improvements could come from scaling to more diverse datasets collected with different devices and careful calibration, as well as expanding coverage to a wider variety of scenes. Employing systematic data curation and incorporating larger models or advanced pre-trained components are also promising directions that could enhance both generalization and overall performance.
5 Conclusion
We presented VoT, an end-to-end method for visual odometry systems based on transformers for direct relative pose prediction. The approach achieves strong results compared to modern state-of-the-art odometry pipelines and large 3D models on the challenging indoor and outdoor datasets. VoT has a flexible architecture easily extensible to various backbones, with competitive results. Importantly, it can generalize to out-of-distribution camera parameters and image data. It exhibits strong scaling behavior, indicating that its performance could be further enhanced with sufficient data and computational resources. Finally, it is considerably faster than alternatives, being more compact than large 3D models, and, unlike other odometry systems, it does not require test-time optimization. We anticipate that future work will further advance end-to-end odometry systems.
Acknowledgements
This work has been financially supported by TomTom, the University of Amsterdam, and the allowance of Top consortia for Knowledge and Innovation (TKIs) from the Netherlands Ministry of Economic Affairs and Climate Policy.